# Decentralized Domain Generalization with Style Sharing: Formal Model and Convergence Analysis

### v3

@@ VI Numerical Experiments
TABLE II: Results on the PACS dataset, for a network of $m=9$ devices over a random geometric graph with radius $r=0.8$.
@@ VI-A Setup
Datasets and Model. Our experiments are based on two widely-used benchmark datasets for DG: (i) PACS [40] which consists of $4$ domains (art painting (A), cartoon (C), photo (P), and sketch (S)) and contains $7$ classes of images (dog, elephant, giraffe, guitar, horse, house, and person); and (ii) VLCS [46] with $4$ domains (Caltech101, LabelMe, PASCAL VOC2007, and SUN09) and $5$ categories of images (bird, car, chair, dog, and person). For both datasets, we use 3 domains as the source domains and reserve the remaining one as the target domain (for evaluating the test performance). By default, we conduct our experiments using the ResNet18 model [39].
@@
Baselines. Since our work is the first to propose a fully decentralized DG algorithm, we select some centralized and federated DG algorithms and implement a decentralized realization of them for comparison. Specifically, we compare StyleDDG against DSGD [43], FedBN [32], $\operatorname{MixStyle}$ [28] and $\operatorname{DSU}$ [29]. For the well-known style-based centralized DG methods, $\operatorname{MixStyle}$ and $\operatorname{DSU}$, we incorporate these algorithms into the local update process at each client.
Setting. Our experiments are done under two sets of networks: (i) $m=3$ devices, and (ii) $m=9$ devices, connected together over a random geometric graph [47]. The choice of 3-9 edge devices aligns with the conventions of experiments found in existing FDG literature [48, 32, 49], thus providing fair comparisons with the baselines. Our implementation of StyleDDG is based on the dassl library developed in [6, 50]. We adopt a constant learning rate $\alpha=0.001$ with a cosine scheduler, and use the batch size of $64$. We distribute the data among the edge devices in a domain-heterogeneous way, where each device receives samples only from a single domain. We measure the test accuracy of each algorithm after $50$ epochs of training. We conduct the experiments based on a cluster of four NVIDIA A100 GPUs with 40GB memory. We run the experiments five times in each setup, and present the mean and standard deviation.
@@ VI-B Results and Discussion
PACS. Table I shows the result with PACS dataset with $m=3$ clients. We make the following observations. First, existing style-based centralized DG methods, i.e., MixStyle and DSU, enhance the performance of both FedBN and DSGD by exposing models to a diversity of styles. However, applying these approaches does not allow the model to sufficiently explore diverse styles, as each device has access to only a limited set of styles or domains. We observe that StyleDDG addresses this limitation through style sharing and exploration in decentralized settings, resulting in a significant improvement in generalization capability.
@@
VLCS. In Table III, we evaluate performance on the VLCS dataset. Although the gains are smaller than those observed on PACS – due to the relatively smaller style gap between domains in VLCS [13]StyleDDG still outperforms on all but one domain.
@@
ResNet50. In Table IV, we explore the DG performance of our algorithm using a larger model, i.e., ResNet50 [39]. Our results show that StyleDDG outperforms all baselines in all but one domain, and is superior to them on average. Also, compared to ResNet18 results from Table I, we see ${\small\sim}7\%$ increase in accuracy of all DG algorithms, which shows that larger models lead to higher-quality style statistics for DG.
Graph Connectivity. In Fig. 3, we vary radius of the random geometric graph in a network of $m=9$ devices and compare the performance of StyleDDG with DSDG for different connectivity levels of the network. We observe that our methodology maintains its superiority across different connectivity levels. Furthermore, the advantage of our method increases with higher graph connectivity, showing that it can better utilize the underlying network topology to achieve DG. In other words, the strength of style sharing in our method becomes more significant if each device is connected to more neighbors, as it can be exposed to a larger set of styles.
@@ VII Conclusion
We proposed StyleDDG, to our knowledge the first decentralized federated domain generalization algorithm. StyleDDG enables edge devices in a decentralized network to train domain-invariant models. We achieved this by forming consensus-based model aggregation and style sharing, and developing a style exploration strategy to expose each device to the new style domains from their neighbors. Further, we provided a novel analytical foundation of well-known domain generalization methods and extended it to our methodology, enabling us to conduct the first convergence analysis among all style-based domain generalization algorithms. Our experiments demonstrated the effectiveness of StyleDDG compared to other domain generalization baselines in terms of model accuracy on unseen target domains.

### v4

@@ VI Numerical Experiments
TABLE II: Results on the PACS dataset, for a network of $m=9$ devices on a random geometric graph with radius $r=0.8$.
@@ VI-A Setup
Datasets and Model. Our experiments are based on two widely-used datasets for benchmarking DG; (i) PACS [li2017deeper] which consists of $4$ domains (art painting (A), cartoon (C), photo (P) and sketch (S)) and contains $7$ classes of images (dog, elephant, giraffe, guitar, horse, house, and person); and (ii) VLCS [torralba2011unbiased] with $4$ domains (Caltech101, LabelMe, PASCAL VOC2007, and SUN09) and $5$ categories of images (bird, car, chair, dog, and person). For both datasets, we use $3$ source domains and hold out the remaining one as the target domain (for evaluating the test performance). By default, we conduct our experiments using the ResNet18 model [he2016deep].
@@
Baselines. Since our work is the first to propose a fully decentralized DG algorithm, we select some centralized and federated DG algorithms and implement a decentralized realization of them for comparison. Specifically, we compare StyleDDG against DSGD [koloskova2020unified], FedBN [lifedbn], $\operatorname{MixStyle}$ [zhou2021domain] and $\operatorname{DSU}$ [li2022uncertainty]. For the well-known style-based centralized DG methods, $\operatorname{MixStyle}$ and $\operatorname{DSU}$, we incorporate these algorithms into the local update process in each client.
Setting. Our experiments are done under two sets of networks: (i) $m=3$ devices and (ii) $m=9$ devices, connected together on a random geometric graph [penrose2003random]. The choice of 33-99 devices aligns with the experiment conventions found in FDG literature [zhang2021federated, lifedbn, nguyen2022fedsr], thus providing fair comparisons with the baselines. Our implementation of StyleDDG is based on the dassl library developed in [zhou2022domain, zhou2021domainadaptive]. We adopt a learning rate $\alpha=0.001$ with a cosine scheduler and use the batch size of $64$. We distribute the data among the edge devices in a domain-heterogeneous way, where each device gets samples only from a single domain. We measure the test accuracy of each algorithm after $50$ training epochs. We carried out the experiments using a cluster of four NVIDIA A100 GPUs with 40GB memory. We run the experiments five times in each setup and present the mean and standard deviation.
@@ VI-B Results and Discussion
PACS. Table I shows the results with the PACS dataset with $m=3$ clients. We make the following observations. First, existing style-based centralized DG methods, i.e., MixStyle and DSU, enhance the performance of both FedBN and DSGD by exposing models to a diversity of styles. However, applying these approaches does not allow the model to sufficiently explore diverse styles in decentralized setups, as each device has access to only a limited set of styles or domains. We observe that StyleDDG addresses this limitation through style sharing and exploration in decentralized settings, resulting in a significant improvement in generalization capability.
@@
VLCS. In Table III, we evaluate performance on the VLCS dataset. Although the gains are smaller than those observed on PACS – due to the relatively smaller style gap between domains in VLCS [park2024stablefdg]StyleDDG still outperforms in all but one domain.
@@
ResNet50. In Table IV, we explore the DG performance of our algorithm using a larger model, i.e., ResNet50 [he2016deep]. Our results show that StyleDDG outperforms all baselines in all but one domain and is superior to them on average. Moreover, compared to the ResNet18 results given in Table I, we see a ${\small\sim}7\%$ increase in accuracy for all DG algorithms, which shows that larger models lead to higher-quality style statistics for DG.
Graph Connectivity. In Fig. 3, we vary the radius of the random geometric graph in a network of $m=9$ devices and compare the performance of StyleDDG with DSDG for different levels of connectivity of the network. We observe that our methodology maintains its superiority across different connectivity levels. Furthermore, the advantage of our method increases with higher graph connectivity, showing that it can better utilize the underlying network topology to achieve DG. In other words, the strength of style sharing in our method becomes more significant if each device is connected to more neighbors, as it can be exposed to a larger set of styles.
@@ VII Conclusion
We proposed StyleDDG, to our knowledge the first decentralized federated domain generalization algorithm. StyleDDG enables edge devices in a decentralized network to train domain-invariant models. We achieved this by forming consensus-based model aggregation, style sharing, and developing a style exploration strategy to expose each device to the new style domains of their neighbors. Furthermore, we provided a novel analytical foundation of well-known domain generalization methods and extended it to our methodology, enabling us to conduct the first convergence analysis among all style-based domain generalization algorithms. Our experiments demonstrated the effectiveness of StyleDDG compared to other domain generalization baselines in terms of model accuracy on unseen target domains.
The authors have provided public access to their code and/or data at https://zenodo.org/records/18204171.
Acknowledgment
This work was supported in part by the U.S. National Science Foundation (NSF) under grant No. CNS-2146171 and ECCS-2512911, the Office of Naval Research (ONR) under grant N00014-22-1-2305, the Defense Advanced Research Projects Agency (DARPA) under grant D22AP00168, and the Air Force Office of Scientific Research (AFOSR) under grant FA9550-24-1-0083.